# 00footnotetext:

### v1

@@ 1 Introduction
Foundation models have revolutionized our world, demonstrating remarkable capabilities in solving grade school math problems, writing creative essays, generating stunning images, and comprehending visual content openai2023gpt4; ChatGPT; ramesh2022hierarchical. One notable example is CLIP (radford2021learning), a vision-language model pretrained on a vast dataset of image-text pairs, which forms the backbone of numerous other foundation models (ramesh2022hierarchical; liu2023visualinstructiontuning). CLIP has achieved unprecedented performance across a wide range of benchmarks spanning many domains—a sharp contrast to models from the ImageNet era, which struggled to generalize from a training domain mostly consisting of natural photographs to stylistically different domains such as ImageNet-Sketch (wang2019learning), ImageNet-R (hendrycks2020many), and DomainNet (peng2019moment).
Domains, while not always rigorously defined, emerge from collecting data from specific sources and conditions. Some domains, like natural images or renditions, are better delineated, allowing the creation of datasets like the ones mentioned above. Out-of-domain (OOD) generalization refers to a model’s ability to perform well on data from domains other than its training domain(s) (wang2022generalizingunseendomainssurvey). In this work, we collectively refer to the domain represented by ImageNet-Sketch, ImageNet-R, DomainNet-Painting, DomainNet-Clipart, DomainNet-Sketch, and DomainNet-Quickdraw as the rendition domain, since it contains images that are renditions of natural objects and scenes. Generalization to the rendition domain (especially OOD) is crucial for aligning models with human perception, as humans can interpret abstract visual renditions, while machines tend to rely heavily on textural cues (hendrycks2020many; geirhos2018imagenet).
@@
In contrast to sample contamination, domain contamination does not care about duplicates of specific datapoints but instead checks whether crucial aspects of a test domain are included in the training domain, e.g., by including images with different content but similar style to test samples. For example, after the correction by mayilvahanan2024, although not duplicates, many other rendition images remained in CLIP’s training set (refer to Tab. LABEL:tab:domain_composition).s from the test domain. This leads us to our central question:
@@ 2 Related Work
We aim to understand the OOD generalization capabilities of CLIP from a data-centric viewpoint. While multi-modal training with rich language captions does seem to contribute to robustness against distribution shifts (xue2024understanding), fang2022data demonstrated that the nature of CLIP’s training distribution (as opposed to its mere size, its specific training objective, or natural language supervision) causes strong performance on various distribution shifts. However, it is unclear what aspects of the data distribution drive the robustness gains. mayilvahanan2024 remove images highly similar to the test sets to show that data contamination and high perceptual similarity between training and test data do not explain generalization performance. While their data pruning technique removes some samples from LAION-400M that lie outside the natural image domain, they do not address domain generalization: They only account for the part of a domain covered by existing test sets and give no guarantee that all images of a given domain were removed. In another line of work, nguyen2022quality discover that a model’s effective robustness (fang2022data; taori2020measuring) on a test set interpolates when training data is compiled from various sources. However, they only consider mixing datasets that each cover multiple domains. In this work, we take their analysis further and show how mixing two data sources from distinct domains interpolates the effective robustness on those domains. Our study’s title is inspired by gulrajani2020search, who studied generalization from multiple distinct source domains. In contrast, we focus on generalization from single or mixed source domains to unseen domains.
@@ 3 Constructing Clean Single-Domain Datasets
To answer our central question—how much of CLIP’s performance on renditions can be explained by domain contamination—we must filter out datapoints from specific domains within web-scale datasets. Similar to how ImageNet is compared to ImageNet-Sketch and ImageNet-R, or how DomainNet-Real is compared to DomainNet-Sketch (Quickdraw, Infograph, Clipart, and Painting), we aim to create clean natural and rendition datasets from LAION by building a domain classifier to distinguish between these domains.
To build a robust domain classifier, we first create a labeled dataset where each class represents a distinct domain. The labeling process is outlined in Sec. 3.1, and we explore different ways to build a domain classifier in Sec. 3.2. Further, in LABEL:sec:testing_domain_classifier, we employ the best-performing classifiers to analyze the composition of different training and test sets and finally use it to subsample LAION-Natural and LAION-Rendition in LABEL:sec:subsampling_datasets.

### v2

@@ 1 Introduction
Foundation models have revolutionized our world, demonstrating remarkable capabilities in solving grade school math problems, writing creative essays, generating stunning images, and comprehending visual content (openai2023gpt4; ChatGPT; ramesh2022hierarchical). One notable example is CLIP (radford2021learning), a vision-language model pretrained on a vast dataset of image-text pairs, which forms the backbone of numerous other foundation models (ramesh2022hierarchical; liu2023visualinstructiontuning). CLIP has achieved unprecedented performance across a wide range of benchmarks spanning many domains—a sharp contrast to models from the ImageNet era, which struggled to generalize from a training domain mostly consisting of natural photographs to stylistically different domains such as ImageNet-Sketch (wang2019learning), ImageNet-R (hendrycks2020many), and DomainNet (peng2019moment).
Domains, while often challenging to quantify in practice (bendavid), emerge from collecting data from specific sources and conditions. Some domains, like natural images or renditions, are better delineated, allowing the creation of datasets like the ones mentioned above. Out-of-domain (OOD) generalization refers to a model’s ability to perform well on data from domains other than its training domain(s) (wang2022generalizingunseendomainssurvey). In this work, we collectively refer to the domain represented by ImageNet-Sketch, ImageNet-R, DomainNet-Painting, DomainNet-Clipart, DomainNet-Sketch, and DomainNet-Quickdraw as the rendition domain, since it contains images that are renditions of natural objects and scenes. Generalization to the rendition domain (especially OOD) is crucial for aligning models with human perception, as humans can interpret abstract visual renditions, while machines tend to rely heavily on textural cues (hendrycks2020many; geirhos2018imagenet).
@@
In contrast to sample contamination, domain contamination does not focus on duplicates of specific datapoints but rather examines whether critical aspects of a test domain are present in the training domain, such as images with different content but similar style to test samples. For example, after the correction by mayilvahanan2024, many other rendition images, while not duplicates, remained in CLIP’s training set (refer to Tab. LABEL:tab:domain_composition). Prior works often assume that CLIP is capable of generalizing OOD (radford2021learning; abbasi2024decipheringrolerepresentationdisentanglement; nguyen2024saftoutofdistributiongeneralizationfinetuning; fang2022data; li2023distillinglargevisionlanguagemodel; shu2023clipoodgeneralizingclipoutofdistributions); however, it remains unclear whether this is truly the case or if its performance is primarily driven by training on images from the test domain. This leads us to our central question:
@@ 2 Related Work
We aim to understand the OOD generalization capabilities of CLIP from a data-centric viewpoint. While multi-modal training with rich language captions does seem to contribute to robustness against distribution shifts (xue2024understanding), fang2022data demonstrated that the nature of CLIP’s training distribution (as opposed to its mere size, its specific training objective, or natural language supervision) causes strong performance on various distribution shifts. However, it is unclear what aspects of the data distribution drive the robustness gains. mayilvahanan2024 remove images highly similar to the test sets to show that data contamination and high perceptual similarity between training and test data do not explain generalization performance. While their data pruning technique removes some samples from LAION-400M that lie outside the natural image domain, they do not address domain generalization: They only account for the part of a domain covered by existing test sets and give no guarantee that all images of a given domain were removed. In another line of work, nguyen2022quality discover that a model’s effective robustness (fang2022data; taori2020measuring) on a test set interpolates when training data is compiled from various sources. However, they only consider mixing datasets that each cover multiple domains. In this work, we take their analysis further and show how mixing two data sources from distinct domains interpolates the effective robustness on those domains. Our study’s title is inspired by gulrajani2020search, who studied generalization from multiple distinct source domains. In contrast, we focus on generalization from single or mixed source domains to unseen domains. Overall, we aim for our work to be a valuable addition to the literature on OOD generalization (liu2023outofdistributiongeneralizationsurvey; koh2021wildsbenchmarkinthewilddistribution; madan2021cnnsgeneralizeoutofdistributioncategoryviewpoint; gulrajani2020search; madan2022what; arjovsky_invariant_2020; arjovsky_out_2021).
@@ 3 Constructing Clean Single-Domain Datasets
To answer our central question—how much of CLIP’s performance on renditions can be explained by domain contamination—we must filter out datapoints from specific domains within web-scale datasets. Similar to how ImageNet is compared to ImageNet-Sketch and ImageNet-R, or how DomainNet-Real is compared to DomainNet-Sketch (Quickdraw, Infograph, Clipart, and Painting), we aim to create clean natural and rendition datasets from LAION by building a domain classifier to distinguish between these domains. To build a robust domain classifier, we first create a labeled dataset where each class represents a distinct domain. The labeling process is outlined in Sec. 3.1, and we explore different ways to build a domain classifier in Sec. 3.2. Further, in LABEL:sec:testing_domain_classifier, we employ the best-performing classifiers to analyze the composition of different training and test sets and finally use it to subsample LAION-Natural and LAION-Rendition in LABEL:sec:subsampling_datasets.